Find an equation of the line that passes through (6,6) and (-6,-6). Write the answer in slope-intercept form, using x as the independent variable.
Answer by alicealc(293)

assume the equation is y = mx + c

pass through (6,6):
6 = 6m + c
pass through (-6,-6):
-6 = -6m + c

use elimination procedure:
6 = 6m + c
-6 = -6m + c
------------- (-)
12 = 12m
m = 12/12 = 1

6 = 6m + c
6 = 6*1 + c
6 = 6 + c
6 - 6 = c
0 = c

so the line equation for the problem is y = x